The radiator in Natalie's car contains 6.3 L of antifreeze and water. This mixture is 30% antifreeze. How much of this mixture should she drain and replace with pure antifreeze so that there will be a mixtuure of 50% antifreeze?
1.89 - .3x + 1x = 3.15

Answer by stanbon(75887)

The radiator in Natalie's car contains 6.3 L of antifreeze and water. This mixture is 30% antifreeze. How much of this mixture should she drain and replace with pure antifreeze so that there will be a mixtuure of 50% antifreeze?
------
Equation:
alcohol - alcohol + alcohol = alcohol
----
0.30*6.3 - 0.30x + 1.00x = 0.50*6.3
--------
30*6.3 - 30x + 100x = 50*6.3
---
70x = 20*6.3
x = (2/7)6.3
x = 2*0.9
x = 1.8 L (amt. to remove and replace)
-------------------
Cheers,
Stan H.